# Upper and Lower Bounds on the Smoothed Complexity of the Simplex Method

### v2

@@ Abstract
The simplex method for linear programming is known to be highly efficient in practice, and understanding its performance from a theoretical perspective is an active research topic. The framework of smoothed analysis, first introduced by Spielman and Teng (JACM ’04) for this purpose, defines the smoothed complexity of solving a linear program with $d$ variables and $n$ constraints as the expected running time when Gaussian noise of variance $\sigma^{2}$ is added to the LP data. We prove that the smoothed complexity of the simplex method is $O(\sigma^{-3/2}d^{13/4}\log^{7/4}n)$, improving the dependence on $1/\sigma$ compared to the previous bound of $O(\sigma^{-2}d^{2}\sqrt{\log n})$. We accomplish this through a new analysis of the shadow bound, key to earlier analyses as well. Illustrating the power of our new approach, we moreover prove a nearly tight upper bound on the smoothed complexity of two-dimensional polygons.
We also establish the first non-trivial lower bound on the smoothed complexity of the simplex method, proving that the shadow vertex simplex method requires at least $\Omega\Big{(}\min\big{(}\sigma^{-1/2}d^{-1/2}\log^{-1/4}d,2^{d}\big{)}\Big{)}$ pivot steps with high probability. A key part of our analysis is a new variation on the extended formulation for the regular $2^{k}$-gon. We end with a numerical experiment that suggests our lower bound could be further improved.
@@ 1 Introduction
Introduced by Dantzig [d47], the simplex method is one of the primary methods for solving linear programs (LP’s) in practice and is an essential component in many software packages for combinatorial optimization. It is a family of local search algorithms which begin by finding a vertex of the set of feasible solutions and iteratively move to a better neighboring vertex along the edges of the feasible polyhedron until an optimal solution is reached. These moves are known as pivot steps. Variants of the simplex method can be differentiated by the choice of pivot rule, which determines which neighbouring vertex is chosen in each iteration, as well as by the method for obtaining the initial vertex. Some well-known pivot rules are the most negative reduced cost rule, the steepest edge rule, and an approximate steepest edge rule known as the devex rule. In theoretical work, the parametric objective rule, also known as the shadow vertex rule, plays an important role.
@@
Empirical evidence suggests that the simplex algorithm typically takes $O(d+n)$ pivot steps, see [sha87, andrei2004complexity, Goldfarb1994]. However, obtaining a rigorous explanation for this excellent performance has proven challenging. In contrast to the practical success of the simplex method, all studied variants are known to have super-polynomial or even exponential worst-case running times. For deterministic variants, many published bad inputs are based on deformed cubes, see [km72, jer73, AC78, GS79, Mur80, gol76] and a unified construction in [jour/cm/AZ98]. For randomized and history-dependent variants, bad inputs have been constructed based on Markov Decision Processes [k92, msw96, hz15, conf/stoc/FHZ11, conf/ipco/Friedmann11, disser2020exponential]. The fastest provable (randomized) simplex algorithm takes $O(2^{\sqrt{d\log n}})$ pivot steps in expectation [k92, msw96, hz15].
@@
Average-case analyses of the simplex method have been performed for a variety of random distributions over linear programs [b82, b87, b99, jour/mapr/Smale83, report/Haimovich83, jour/mapr/Megiddo86, jour/jc/AKS87, jour/mapr/Todd86, jour/jacm/AM85]. While insightful, the results from average-case analyses might not be fully realistic due to the fact that “random” linear programs tend to have certain properties that “typical” linear programs do not.
@@
To better explain why simplex algorithm performs well in practice, while avoiding some of the pitfalls of average-case analysis, Spielman and Teng [ST04] introduced the smoothed complexity framework. For any base LP data $\bar{A}\in\mathbb{R}^{n\times d},\bar{b}\in\mathbb{R}^{n},c\in\mathbb{R}^{d}\backslash\{0\}$ where the rows of $(\bar{A},\bar{b})$ are normalized to have $\ell_{2}$ norm at most $1$, they consider the smoothed LP obtained by adding independent Gaussian perturbations to the constraints:
$\max_{x\in\mathbb{R}^{d}}c^{\top}x\quad\text{subject to}\quad(\bar{A}+\hat{A})x\leq(\bar{b}+\hat{b}).$
@@
The entries of $\hat{A}$ and $\hat{b}$ are i.i.d. Gaussian random variables with mean 00 and variance $\sigma^{2}$. The smoothed complexity of a simplex algorithm $\mathcal{A}$ is defined to be the maximum (over $\bar{A},\bar{b},c$) expected number of pivot steps the algorithm takes to solve the smoothed LP, i.e.,
@@
Here $T_{\mathcal{A}}(A,b,c)$ is the number of pivot steps that the algorithm $\mathcal{A}$ takes to solve the linear program $\max_{x\in\mathbb{R}^{d}}\{c^{\top}x:~{}Ax\leq b\}$. We may note that if $\sigma\to\infty$ then $\mathcal{SC}_{\mathcal{A},n,d,\sigma}$ approaches the average-case complexity of $\mathcal{A}$ on independent Gaussian distributed input data. In constrast, if $\sigma\to 0$ then $\mathcal{SC}_{\mathcal{A},n,d,\sigma}$ will approach the worst-case complexity of $\mathcal{A}$. As a result, most interest has been directed at understanding the dependence on $\sigma$ in the regime where $\sigma\geq 2^{-\Omega(d)}$ but $\sigma\leq 1/\operatorname{poly}(d)$.
@@
An algorithm is said to have polynomial smoothed complexity if under the perturbation of constraints, it has expected running time $\operatorname{poly}(n,d,\sigma^{-1})$, and [ST04] proved that the smoothed complexity of a specific simplex method based on the shadow vertex simplex method (which we will describe next) is at most $O(d^{55}n^{86}\sigma^{-30}+d^{70}n^{86})$. The best bound available in the literature is $O(\sigma^{-2}d^{2}\sqrt{\log n})$ pivot steps due to [DH18], assuming $\sigma\leq 1/\sqrt{d\log n}$. We note that assuming an upper bound on $\sigma$ can be done without loss of generality; its influence can be captured as an additive term in the upper bound that does not depend on $\sigma$.
This work improves the dependence on $\sigma$ of the smoothed complexity, obtaining an upper bound of $O(\sigma^{-3/2}d^{13/4}\log^{7/4}n)$ for $\sigma\leq 1/d\sqrt{\log n}$. As a second contribution, we prove the first non-trivial lower bound on the smoothed complexity of a simplex method, finding that the shadow vertex simplex method requires $\Omega(\min(\frac{1}{\sqrt{\sigma d\sqrt{\log n}}},2^{d}))$ pivot steps.
@@
One of the most extensively studied simplex algorithms in theory is the shadow vertex simplex algorithm [gas55, b82]. Given an LP
@@
for $A\in\mathbb{R}^{n\times d},b\in\mathbb{R}^{n},c\in\mathbb{R}^{d}$, let $P=\{x\in\mathbb{R}^{d}:Ax\leq b\}$ denote the feasible polyhedron of the linear program. The algorithm starts from an initial vertex $x_{0}\in P$ that optimizes an initial objective $c_{0}$ [^1] There are many standard methods of finding such initialization with at most multiplicative $O(d)$ overhead in running time, so we can assume that both $x_{0}$ and $c_{0}$ are already given. See the discussion in [DH18]. During the execution, it maintains an intermediate objective $c_{\lambda}=\lambda c+(1-\lambda)c_{0}$ and a vertex that optimizes $c_{\lambda}$. Thus by slowly increasing $\lambda$ from 00 to $1$ during different pivot steps, the temporary objective gradually changes from $c_{0}$ to $c$, revealing the desired solution at the end. Since each pivot step requires $\operatorname{poly}(d,n)$ computational work, theoretical analysis has focused on analyzing the number of pivot steps.
The algorithm is called shadow vertex simplex method because, when performing orthogonal projection of the feasible set onto the two-dimensional linear subspace $W=\mathrm{span}(c_{0},c)$, the vertices visited by the algorithm project onto the boundary of the projection (“shadow”) $\pi_{W}(P)$. Assuming certain non-degeneracy conditions, which will hold with probability $1$ for the distributions we consider, this projection gives an injective map from iterations of the method to vertices of the shadow, meaning that we can upper bound the number of pivot steps in the algorithm by the number of vertices of the shadow polygon. This characterization makes the shadow vertex simplex method ideally suited for probabilistic analysis.
To analyze the “shadow size”, the number of vertices of the shadow polygon, we follow earlier work of [ver09] and reduce to the case that $b=\mathbf{1}_{n}$. In this case, well-established principles of polyhedral duality show that we can bound the number of vertices of a convex polygon by the number of edges of its dual polygon:
@@
The following upper bound we take from [DH18], which states that the analysis of [ver09] can be strengthened to obtain the claimed bound. This upper bound should be understood as stating that there exists a shadow vertex rule based simplex algorithm which satisfies that smoothed complexity bound. The lower bound is due to [b87] and shows that the shadow vertex simplex rule, with two adversarially specified objectives $c_{0},c$, can be made to follow paths of this length.
Theorem 1 (Smoothed Complexity of Shadow Vertex Simplex Algorithm).
@@
With this reduction, analyzing the smoothed complexity of the simplex method comes down to bounding the smoothed shadow size $\mathcal{S}(n,d,\sigma)$. As such, that will be the focus of the remainder of this paper.
@@ 1.1 Our Results
The previous best shadow bound is due to [DH18], who prove that $\mathcal{S}(n,d,\sigma)\leq O(d^{2}\sqrt{\log n}\sigma^{-2})$. Our first main result strengthen this result for small values of $\sigma$.
Theorem 2.
@@
$\mathcal{S}(n,d,\sigma)=O\left(\sigma^{-3/2}d^{13/4}\log^{7/4}n\right).$
@@
Second, we prove the first non-trival lower bound on the smoothed shadow size, establishing that $\mathcal{S}(4d-13,d,\sigma)\geq\Omega(\min(\frac{1}{\sqrt{\sigma d\sqrt{\log d}}},2^{d}))$ for $d>5$. This lower bound is proven by constructing a polyhedron $P=\{x\in\mathbb{R}^{d}:Ax\leq\mathbf{1}_{n}\}$ and a two-dimensional subspace $W$ such that for any small perturbation of $A$, the new polyhedron $P$ projected onto $W$ will have many vertices. The construction is based on an extended formulation similar to those first constructed by [BN01, glineur].
Theorem 3.
@@
[b87] $\Theta(d^{3/2}\sqrt{\log n})$ Average- case, Gaussian
[ST04] $O(\sigma^{-6}d^{3}n+d^{6}n\log^{3}n)$ Smooth
[ds05] $O(\sigma^{-2}dn^{2}\log n+d^{2}n^{2}\log^{2}n)$ Smooth
[ver09] $O(\sigma^{-4}d^{3}+d^{5}\log^{2}n)$ Smooth
[DH18] $O(\sigma^{-2}d^{2}\sqrt{\log n}+d^{3}\log^{1.5}n)$ Smooth
This paper $O(\sigma^{-3/2}d^{13/4}\log^{7/4}n+d^{19/4}\log^{13/4}n)$ Smooth
@@
[Mur80, report/Goldfarb83] $2^{d}$ Worst
Table 1: Bounds of expected number of pivots in previous literature, assuming $d\geq 3$. Logarithmic factors are simplified. The lower bound of [b87] holds in the smoothed model as well.
@@
To better understand the smoothed complexity of the intersection polygon $\operatorname{conv}(a_{1},\dots,a_{n})\cap W$, we also analyze its two-dimensional analogue introduced by [damerow2004extreme]. Taking $\bar{a}_{1},\dots,\bar{a}_{n}\in\mathbb{R}^{2}$, each satisfying $\|a_{i}\|_{2}\leq 1$, we are interested in the number of edges of the smoothed polygon $\operatorname{conv}(\bar{a}_{1}+\hat{a}_{1},\dots,\bar{a}_{n}+\hat{a}_{n})$, where $\hat{a}_{1},\dots,\hat{a}_{n}\sim N(0,\sigma^{2})$ are independent. The previous best upper bound on the smoothed complexity of this polygon is $O(\sigma^{-1}+\sqrt{\log n})$, due to [bwcachapter]. Their analysis is based on an adaptation of the shadow bound by [DH18]. In Section 4 we improve this upper bound, obtaining the following theorem.
Theorem 4 (Two-Dimensional Upper Bound).
@@
To confirm that the above upper bound is stronger than that of [bwcachapter], one may use the AM-GM inequality to verify that $2\frac{\sqrt[4]{\log n}}{\sqrt{\sigma}}\leq\sqrt{\log n}+\sigma^{-1}$. Combined with the trivial upper bound of $n$ vertices, our bound nearly matches the lower bound of $\Omega(\min(\sqrt{\log n}+\frac{\sqrt[4]{\log(n\sqrt{\sigma})}}{\sqrt{\sigma}},n))$ proven by [DGGT16]. A full overview of previous results on the smoothed complexity of the two-dimensional convex hull can be found in Table 2.
@@
[damerow2004extreme] $O(\log(n)^{2}+\sigma^{-2}\log n)$
[thesis/Schnalzger14] $O(\log n+\sigma^{-2})$
[DGGT16] $O(\sqrt{\log n}+\sigma^{-1}\sqrt{\log n})$
[bwcachapter] $O(\sqrt{\log n}+\sigma^{-1})$
@@
[DGGT16] $\Omega(\min(\sqrt{\log n}+\frac{\sqrt[4]{\log(n\sqrt{\sigma})}}{\sqrt{\sigma}},n))$

### v3

@@ Abstract
The simplex method for linear programming is known to be highly efficient in practice, and understanding its performance from a theoretical perspective is an active research topic. The framework of smoothed analysis, first introduced by Spielman and Teng (JACM ’04) for this purpose, defines the smoothed complexity of solving a linear program with $d$ variables and $n$ constraints as the expected running time when Gaussian noise of variance $\sigma^{2}$ is added to the LP data. We prove that the smoothed complexity of the simplex method is $O(\sigma^{-3/2}d^{13/4}\log^{5/4}n)$, improving the dependence on $1/\sigma$ compared to the previous bound of $O(\sigma^{-2}d^{2}\sqrt{\log n})$. We accomplish this through a new analysis of the shadow bound, key to earlier analyses as well. Illustrating the power of our new approach, we moreover prove a nearly tight upper bound on the smoothed complexity of two-dimensional polygons.
We also establish the first non-trivial lower bound on the smoothed complexity of the simplex method, proving that the shadow vertex simplex method requires, with a given auxiliary objective, at least $\Omega\Big(\min\big(\sigma^{-1/2}d^{-1/2}\log^{-1/4}d,2^{d}\big)\Big)$ pivot steps with high probability. A key part of our analysis is a new variation on the extended formulation for the regular $2^{k}$-gon. We end with a numerical experiment that suggests our lower bound could be further improved.
@@ 1 Introduction
Introduced by Dantzig [20], the simplex method is one of the primary methods for solving linear programs (LP’s) in practice and is an essential component in many software packages for combinatorial optimization. It is a family of local search algorithms which begin by finding a vertex of the set of feasible solutions and iteratively move to a better neighboring vertex along the edges of the feasible polyhedron until an optimal solution is reached. These moves are known as pivot steps. Variants of the simplex method can be differentiated by the choice of pivot rule, which determines which neighboring vertex is chosen in each iteration, as well as by the method for obtaining the initial vertex. Some well-known pivot rules are the most negative reduced cost rule, the steepest edge rule, and an approximate steepest edge rule known as the Devex rule. In theoretical work, the parametric objective rule, also known as the shadow vertex rule, plays an important role.
@@
Empirical evidence suggests that the simplex algorithm typically takes $O(d+n)$ pivot steps, see [47, 5, 29]. However, obtaining a rigorous explanation for this excellent performance has proven challenging. In contrast to the practical success of the simplex method, all studied variants are known to have super-polynomial or even exponential worst-case running times. For deterministic variants, many published bad inputs are based on deformed cubes, see [40, 36, 6, 32, 43, 30] and a unified construction in [4]. For randomized and history-dependent variants, bad inputs have been constructed based on Markov Decision Processes [38, 41, 25, 24, 34, 23]. The fastest provable (randomized) simplex algorithm takes $O(2^{\sqrt{d\log n}})$ pivot steps in expectation [38, 41, 34].
@@
Average-case analyses of the simplex method have been performed for a variety of random distributions over linear programs [14, 13, 12, 48, 33, 42, 1, 50, 2]. While insightful, the results from average-case analyses might not be fully realistic due to the fact that “random” linear programs tend to have certain properties that “typical” linear programs do not.
@@
To better explain why the simplex algorithm performs well in practice, while avoiding some of the pitfalls of average-case analysis, Spielman and Teng [49] introduced the smoothed complexity framework. For any base LP data $\bar{A}\in\mathbb{R}^{n\times d},\bar{b}\in\mathbb{R}^{n},c\in\mathbb{R}^{d}\backslash\{0\}$ where the rows of $(\bar{A},\bar{b})$ are normalized to have $\ell_{2}$ norm at most $1$, they consider the smoothed LP obtained by adding independent Gaussian perturbations to the constraints:
[…]
@@
The entries of $\hat{A}$ and $\hat{b}$ are i.i.d. Gaussian random variables with mean 0 and variance $\sigma^{2}$. The smoothed complexity of a simplex algorithm $\mathcal{A}$ is defined to be the maximum (over $\bar{A},\bar{b},c$) expected number of pivot steps the algorithm takes to solve the smoothed LP, i.e.,
@@
Here $T_{\mathcal{A}}(A,b,c)$ is the number of pivot steps that the algorithm $\mathcal{A}$ takes to solve the linear program $\max_{x\in\mathbb{R}^{d}}\{c^{\top}x:~Ax\leq b\}$. We may note that if $\sigma\to\infty$ then $\mathcal{SC}_{\mathcal{A},n,d,\sigma}$ approaches the average-case complexity of $\mathcal{A}$ on independent Gaussian distributed input data. In contrast, if $\sigma\to 0$ then $\mathcal{SC}_{\mathcal{A},n,d,\sigma}$ will approach the worst-case complexity of $\mathcal{A}$. As a result, most interest has been directed at understanding the dependence on $\sigma$ in the regime where $\sigma\geq 2^{-\Omega(d)}$ but $\sigma\leq 1/\operatorname{poly}(d)$.
@@
An algorithm is said to have polynomial smoothed complexity if under the perturbation of constraints, it has expected running time $\operatorname{poly}(n,d,\sigma^{-1})$, and [49] proved that the smoothed complexity of a specific simplex method based on the shadow vertex simplex method (which we will describe next) is at most $O(d^{55}n^{86}\sigma^{-30}+d^{70}n^{86})$. The best bound available in the literature is $O(\sigma^{-2}d^{2}\sqrt{\log n})$ pivot steps due to [17], assuming $\sigma\leq 1/\sqrt{d\log n}$. We note that assuming an upper bound on $\sigma$ can be done without loss of generality; its influence can be captured as an additive term in the upper bound that does not depend on $\sigma$.
This work improves the dependence on $\sigma$ of the smoothed complexity, obtaining an upper bound of $O(\sigma^{-3/2}d^{13/4}\log^{5/4}n)$ for $\sigma\leq 1/d\sqrt{\log n}$. As a second contribution, we prove the first non-trivial lower bound on the smoothed complexity of a simplex method, finding that the shadow vertex simplex method requires $\Omega(\min(\frac{1}{\sqrt{\sigma d\sqrt{\log n}}},2^{d}))$ pivot steps.
@@
One of the most extensively studied simplex algorithms in theory is the shadow vertex simplex algorithm [27, 14]. Given an LP
@@
for $A\in\mathbb{R}^{n\times d},b\in\mathbb{R}^{n},c\in\mathbb{R}^{d}$, let $P=\{x\in\mathbb{R}^{d}:Ax\leq b\}$ denote the feasible polyhedron of the linear program. The algorithm starts from an initial vertex $x_{0}\in P$ that optimizes an initial objective $c_{0}$ [^1] There are many standard methods of finding such initialization with at most multiplicative $O(d)$ overhead in running time, so we can assume that both $x_{0}$ and $c_{0}$ are already given. See the discussion in [17].. During the execution, it maintains an intermediate objective $c_{\lambda}=\lambda c+(1-\lambda)c_{0}$ and a vertex that optimizes $c_{\lambda}$. Thus by slowly increasing $\lambda$ from 0 to $1$ during different pivot steps, the temporary objective gradually changes from $c_{0}$ to $c$, revealing the desired solution at the end. Since each pivot step requires $\operatorname{poly}(d,n)$ computational work, theoretical analysis has focused on analysing the number of pivot steps.
The algorithm is called the shadow vertex simplex method because, when performing orthogonal projection of the feasible set onto the two-dimensional linear subspace $W=\mathrm{span}(c_{0},c)$, the vertices visited by the algorithm project onto the boundary of the projection (“shadow”) $\pi_{W}(P)$. Assuming certain non-degeneracy conditions, which will hold with probability $1$ for the distributions we consider, this projection gives an injective map from iterations of the method to vertices of the shadow, meaning that we can upper bound the number of pivot steps in the algorithm by the number of vertices of the shadow polygon. This characterization makes the shadow vertex simplex method ideally suited for probabilistic analysis.
To analyse the “shadow size”, the number of vertices of the shadow polygon, we follow earlier work of [52] and reduce to the case that $b=\mathbf{1}_{n}$. In this case, well-established principles of polyhedral duality show that we can bound the number of vertices of a convex polygon by the number of edges of its dual polygon:
@@
The following upper bound we take from [17], which states that the analysis of [52] can be strengthened to obtain the claimed bound. This upper bound should be understood as stating that there exists a shadow vertex rule based simplex algorithm which satisfies that smoothed complexity bound. The lower bound is due to [13] and shows that the shadow vertex simplex rule, with two adversarially specified objectives $c_{0},c$, can be made to follow paths of this length.
Theorem 1.1 (Smoothed Complexity of Shadow Vertex Simplex Algorithm).
@@
With this reduction, analysing the smoothed complexity of the simplex method comes down to bounding the smoothed shadow size $\mathcal{S}(n,d,\sigma)$. As such, that will be the focus of the remainder of this paper.
@@ 1.1 Our Results
The previous best shadow bound is due to [17], who prove that $\mathcal{S}(n,d,\sigma)\leq O(d^{2}\sqrt{\log n}\sigma^{-2})$. Our first main result strengthens this result for small values of $\sigma$.
Theorem 1.2.
@@
$\mathcal{S}(n,d,\sigma)=O\left(\sigma^{-3/2}d^{13/4}\log^{5/4}n\right).$
@@
Second, we prove the first non-trivial lower bound on the smoothed shadow size, establishing that $\mathcal{S}(4d-13,d,\sigma)\geq\Omega(\min(\frac{1}{\sqrt{\sigma d\sqrt{\log d}}},2^{d}))$ for $d>5$. This lower bound is proven by constructing a polyhedron $P=\{x\in\mathbb{R}^{d}:Ax\leq\mathbf{1}_{n}\}$ and a two-dimensional subspace $W$ such that for any small perturbation of $A$, the new polyhedron $P$ projected onto $W$ will have many vertices. The construction is based on an extended formulation similar to those first constructed by [8, 28].
Theorem 1.3.
@@
[13] $\Theta(d^{3/2}\sqrt{\log n})$
Average-
case,
Gaussian
[49] $O(\sigma^{-6}d^{3}n+d^{6}n\log^{3}n)$ Smooth
[21] $O(\sigma^{-2}dn^{2}\log n+d^{2}n^{2}\log^{2}n)$ Smooth
[52] $O(\sigma^{-4}d^{3}+d^{5}\log^{2}n)$ Smooth
[17] $O(\sigma^{-2}d^{2}\sqrt{\log n}+d^{3}\log^{1.5}n)$ Smooth
This paper $O(\sigma^{-3/2}d^{13/4}\log^{5/4}n+d^{19/4}\log^{2}n)$ Smooth
@@
[53, 43, 31] $2^{d}$ Worst
Table 1: Bounds of expected number of pivots in previous literature, assuming $d\geq 3$. Logarithmic factors are simplified. The lower bound of [13] holds in the smoothed model as well.
@@
To better understand the smoothed complexity of the intersection polygon $\operatorname{conv}(a_{1},\dots,a_{n})\cap W$, we also analyse its two-dimensional analogue introduced by [19]. Taking $\bar{a}_{1},\dots,\bar{a}_{n}\in\mathbb{R}^{2}$, each satisfying $\|\bar{a}_{i}\|_{2}\leq 1$, we are interested in the number of edges of the smoothed polygon $\operatorname{conv}(\bar{a}_{1}+\hat{a}_{1},\dots,\bar{a}_{n}+\hat{a}_{n})$, where $\hat{a}_{1},\dots,\hat{a}_{n}\sim N(0,\sigma^{2})$ are independent. The previous best upper bound on the smoothed complexity of this polygon is $O(\sigma^{-1}+\sqrt{\log n})$, due to [18]. Their analysis is based on an adaptation of the shadow bound by [17]. In Section 4 we improve this upper bound, obtaining the following theorem.
Theorem 1.4 (Two-Dimensional Upper Bound).
@@
To confirm that the above upper bound is stronger than that of [18], one may use the AM-GM inequality to verify that $2\frac{\sqrt[4]{\log n}}{\sqrt{\sigma}}\leq\sqrt{\log n}+\sigma^{-1}$. Combined with the trivial upper bound of $n$ vertices, our bound nearly matches the lower bound of $\Omega(\min(\sqrt{\log n}+\frac{\sqrt[4]{\log(n\sqrt{\sigma})}}{\sqrt{\sigma}},n))$ proven by [22]. A full overview of previous results on the smoothed complexity of the two-dimensional convex hull can be found in Table 2.
@@
[19] $O(\log(n)^{2}+\sigma^{-2}\log n)$
[45] $O(\log n+\sigma^{-2})$
[22] $O(\sqrt{\log n}+\sigma^{-1}\sqrt{\log n})$
[18] $O(\sqrt{\log n}+\sigma^{-1})$
@@
[22] $\Omega(\min(\sqrt{\log n}+\frac{\sqrt[4]{\log(n\sqrt{\sigma})}}{\sqrt{\sigma}},n))$